OpenStudy (anonymous):

Solve the following quadratic equation by completing the square: x2 + 5x - 3 = 0 A. (-5+√37)/2, (-5 -√37)/2 B. (5+√37)/2, (5 -√37)/2 C. (-5+√13)/2, (-5 -√13)/2 D. No real solution

OpenStudy (anonymous):

another i need help is Solve for x by completing the square: 4x2 + 6x - 6 = 0 A. (-3+√33)/4, (-3-√33)/4 B. (3+√33)/4, (3-√33)/4 C. 3, -2 D. 1, 6

OpenStudy (anonymous):

completing: \[x^{2} + bx = (x + (b/2))^{2} - (b/2)^{2}\]

OpenStudy (anonymous):

gotta go sorry lol

OpenStudy (anonymous):

its okay thanks :D

OpenStudy (primeralph):

first off, divide all by a = 4

OpenStudy (anonymous):

\[x^2+ 5x- 3 =0 \\\\ x^2+ 5x= 3 \\ x^2+ 5x +(\frac{ 5 }{2})^2 = 3 +(\frac{5 }{2})^2\\\\ x^2+ 5x + (\frac{ 25 }{4})= 3 + (\frac{ 25 }{4})\\\\ ( x + \frac{ 5 }{2} )^2 = \frac{37 }{4} \\\\ x = - \frac{ 5 }{2} \pm \frac{ \sqrt{ 37 }}{ 2} \\\\x = - \frac{ 5 - \sqrt{ 37 }}{2} , - \frac{ 5 + \sqrt{ 37 }}{2} \]